OpenStudy (dmezzullo):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $1,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain ho

OpenStudy (dmezzullo):

help

OpenStudy (dmezzullo):

?

OpenStudy (anonymous):

As it is simple interest, use the formula Interest = Principal * Rate * Time, then you'll get the interest value. Add that to your original value, and calculate the monthly average~

OpenStudy (dmezzullo):

60p+1,000=16,000 -1,000=-1000 60p=15,000 $250 dollars for option 1. 16,000 at 2percent finance 15,000x0.02x6 $1920 dollars interest. Option 2 is the lower amount.

OpenStudy (dmezzullo):

is tht right?

OpenStudy (anonymous):

Yh, I got that too~

OpenStudy (dmezzullo):

sweeet